OpenStudy (anonymous):

y varies directly as x, and y = -1\4 when x = 8. a. Find constant of variation and b. Find y when x = 12. Please explain this to me as well as help me solve this

jimthompson5910 (jim_thompson5910):

y varies directly as x means that y = kx where k is some fixed number

jimthompson5910 (jim_thompson5910):

we are told that "y = -1/4 when x = 8" so y = kx -1/4 = k*8 solve for k

OpenStudy (anonymous):

So then, divide -1\4 by 8?

jimthompson5910 (jim_thompson5910):

to get k = ???

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

yes but what do you get for k when you do that

OpenStudy (anonymous):

-0.031

jimthompson5910 (jim_thompson5910):

keep it as a fraction

OpenStudy (anonymous):

so -31\100?

jimthompson5910 (jim_thompson5910):

-1/4 = k*8 (-1/4)/8 = k k = (-1/4)/8 k = (-1/4)*(1/8) k = ??

OpenStudy (anonymous):

um...-1 over 32?

OpenStudy (anonymous):

That's what my textbook says the answer is

jimthompson5910 (jim_thompson5910):

that is correct, k = -1/32

OpenStudy (anonymous):

But when you take -1\4 divided by 1\8 you have to multiply -1\4 by 8 over 1 which ends up being 2, so how did they get -1\32?

jimthompson5910 (jim_thompson5910):

no, when you divide -1/4 by 8, you are doing this (-1/4)/8 (-1/4)*(1/8) (-1*1)/(4*8) -1/32

OpenStudy (anonymous):

Or do you not multiply by the reciprocal of 1\8?

OpenStudy (anonymous):

Oh I think I get it

jimthompson5910 (jim_thompson5910):

because remember, when dividing fractions, you flip the second fraction and multiply

OpenStudy (anonymous):

So for the second half of the question, you take -1\4 and multiply it by the opposite of 12?

jimthompson5910 (jim_thompson5910):

what was k again

jimthompson5910 (jim_thompson5910):

remember the general formula is y = kx

OpenStudy (anonymous):

Oh wait...k is -1\32, which means you take -1\4 = -1\32 times 12

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

y = kx y = (-1/32)*x ... replace k with -1/32 now you plug in x = 12 to get y = (-1/32)*12 y = ??

OpenStudy (anonymous):

-3\8

jimthompson5910 (jim_thompson5910):

yep, y = -3/8 btw it's not a big deal, but your fraction slashes are going the wrong way (just wanted to let you know)

OpenStudy (anonymous):

So if s varies directly as the square of t and s and t both equal 10, would you use the inverse variation to solve it?

jimthompson5910 (jim_thompson5910):

no it says directly

jimthompson5910 (jim_thompson5910):

not inversely

OpenStudy (anonymous):

So then, would be s = kt^2 which would turn it into 10 = k (10)^2 ?

jimthompson5910 (jim_thompson5910):

good, now solve for k

jimthompson5910 (jim_thompson5910):

keep it as a fraction

OpenStudy (anonymous):

1\10 ?

jimthompson5910 (jim_thompson5910):

1/10 yes, good

OpenStudy (anonymous):

When I make 5 equal to t, I keep getting 2\5 when it should be 5\2...what am I doing wrong?

jimthompson5910 (jim_thompson5910):

s = kt^2 s = (1/10)t^2 s = (1/10)(5)^2 s = (1/10)*25 s = 25/10 s = 5/2

jimthompson5910 (jim_thompson5910):

not sure, but you somehow flipped things somewhere

OpenStudy (anonymous):

I just did 10\1 instead of 1\10

jimthompson5910 (jim_thompson5910):

ah, you have to keep k the same